OpenStudy (anonymous):

Which of the following points lies on the circle whose center is at the origin and whose radius is 5? (1,-2)?

OpenStudy (anonymous):

For center at origin the equation of circle simply becomes: \[\large x^2 + y^2 = r^2\] Or: \[\large x^2 + y^2 = 25\] Now for values of x and y you are getting 25 ??

OpenStudy (anonymous):

\[(\sqrt{5},\sqrt{5})\]

OpenStudy (anonymous):

\[(\sqrt{5})^2 + (\sqrt{5})^2 = 5 + 5 = 10 \ne 25\]

OpenStudy (anonymous):

hmm (-3,4).

OpenStudy (anonymous):

Yes....

OpenStudy (anonymous):

Are you in a test or quiz @helpme9283

OpenStudy (anonymous):

no, i'm studying my own with this math problem books